OpenStudy (anonymous):

WILL MEDAL NEED ANSWER FAST n > - 120/p

OpenStudy (anonymous):

@SWAGSTER23780

OpenStudy (anonymous):

sorry, i do not now how to solve this ='(

OpenStudy (anonymous):

ok do you know anyone who could help me

OpenStudy (anonymous):

120+n-2 whats n equal to? 5 5-120=-115+160=45

OpenStudy (anonymous):

so is 45 the answer

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

what about this Solve 5m − 3h = 45 for h. Show your work

OpenStudy (anonymous):

@masterlot168

OpenStudy (anonymous):

@Xyeana Solve 5m − 3h = 45 for h. Show your work

OpenStudy (anonymous):

5m − 3h = 45 subtract 5m on both sides 5m-5m-3h=45-5m -3h=45-5m divided -3 on both sides -3h/-3=(45-5m)/-3 h=-45/3+5m/3 h=-15+5m/3